# SONAR: A Synthetic AI-Audio Detection Framework and Benchmark

### v3

@@ 4.2.1 How well can detection models generalize across datasets?
Speech foundation models exhibit stronger generalizability. As shown in Table  2, when evaluated on the test set of Wavefake, all models demonstrate near-perfect performance across the three metrics. This can be attributed to the similarity between the test set and the training data. However, when tested on the LibriSeVoc and In-the-wild datasets, models such as LFCC-LCNN, Spec.+ResNet, RawNet2, RawGATST, and AASIST struggle to generalize effectively. This performance gap indicates significant overfitting to the training data, despite these models being specifically designed for audio deepfake detection tasks. In contrast, speech foundation models consistently display stronger generalizability. Notably, Wave2Vec2BERT achieves the highest generalizability, which may be attributed to its large-scale and diverse pretraining data. Pretrained on 4.5 million hours of unlabeled audio in more than 143 languages, Wave2Vec2BERT benefits from both scale and diversity. This suggests that a well-designed self-supervised model trained on diverse speech data can extract general and discriminative features, making it more applicable across different datasets for audio deepfake detection. It is important to note that CLAP, unlike other speech foundation models, does not generalize well across datasets. This is likely due to its primary focus on environmental audio data during pretraining, resulting in the extraction of irrelevant features for speech audio. This observation underscores that not all foundation models are equally suited for audio deepfake detection tasks.
Generalizability may increase with model size. In Table 2, it can be observed that Whisper-large always outperforms Whisper-small across all three datasets. In particular, on the LibriSeVoc test set, Whisper-large achieves accuracy, AUROC, and EER of 0.9572, 0.9901, 4.279%, respectively, which improves by 2.27%, 0.64%, and 2.272%, than that of Whisper-small. This trend is more evident in the In-the-wild dataset, which is closer to real-world scenarios since this dataset consists of speech data sourced from the internet. Specifically, Whisper-large achieves accuracy, AUROC, and EER of 0.8848, 0.9552, and 11.518%, respectively, which improves by 6.381%, 5.27%, and 6.381%, than that of Whisper-small. Further investigation will be made in Section 4.2.3
4.2.2 Results on SONAR dataset
@@
Table 3: Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
@@
4.2.3 Can generalizability increase with model size?
Building on the observation that Whisper-large consistently outperforms Whisper-small, we extend our analysis with controlled experiments on the entire Whisper model family. Specifically, the Whisper family comprises five different model sizes: Whisper-tiny, Whisper-base, Whisper-small, Whisper-medium, and Whisper-large. Table 4 presents the number of model parameters of them. Specifically, each model is fine-tuned on the Wavefake training dataset using the same hyperparameters. Our results show that as model size increases, the generalizability of the models improves as well.
Table 4: Whisper model sizes.
@@ 4.2.3 Can generalizability increase with model size?
Table 5: Generalization across existing audio deepfake datasets. All Whisper models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.
@@
Table 6: Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
@@
4.2.4 On the effectiveness and efficiency of few-shot fine-tuning to improve generalization
Despite the challenges in generalizing across different datasets, we investigate whether there exist efficient solutions that can enhance models’ detection performance on those challenging subsets from SONAR dataset. To this end, we conduct a case study on Wave2Vec2BERT and HuBERT, as these models perform relatively poorly on the OpenAI and SeedTTS datasets but demonstrate competitive performance on other subsets. Specifically, we generate 100 additional fake audio samples using the OpenAI TTS API and randomly select another 100 fake audio samples from the SeedTTS test set for few-shot fine-tuning. Our study yields several interesting findings.
Figures 2(a) and 2(b) present the results of fine-tuning Wave2Vec2BERT and HuBERT using varying numbers of samples from OpenAI. Before fine-tuning, Wave2Vec2BERT and HuBERT only achieve accuracies of 0.7833 and 0.5658, respectively. Notably, with only 10 shots of fake speech data, Wave2Vec2BERT reaches an accuracy of approximately 0.97, while HuBERT’s accuracy increases significantly to approximately 0.85. Importantly, the models’ generalization to other datasets remains unchanged, demonstrating the effectiveness and efficiency of few-shot fine-tuning. However, as the number of fine-tuning samples increases, HuBERT’s test accuracy on the WaveFake test set shows a declining trend, which is also observed for Wave2Vec2BERT.
@@ 4.2.4 On the effectiveness and efficiency of few-shot fine-tuning to improve generalization
Figure 2: Performance of few-shot fine-tuning for Wave2Vec2BERT and HuBERT with a varying number of few-shot audio samples from OpenAI and Seed-TTS, respectively. (a) Fine-tune Wave2Vec2BERT on OpenAI. (b) Fine-tune HuBERT on OpenAI; (c) Fine-tune Wave2Vec2BERT on Seed-TTS; and (d) Fine-tune HuBERT on Seed-TTS.
5 Discussion

### v4

@@ 4.2.1. How well can detection models generalize across datasets?
Speech foundation models exhibit stronger generalizability. As shown in Table  3, when evaluated on the test set of Wavefake, all models demonstrate near-perfect performance across the three metrics. This can be attributed to the similarity between the test set and the training data. However, when tested on the LibriSeVoc and In-the-wild datasets, models such as LFCC-LCNN, Spec.+ResNet, RawNet2, RawGATST, and AASIST struggle to generalize effectively. This performance gap indicates significant overfitting to the training data, despite these models being specifically designed for audio deepfake detection tasks. In contrast, speech foundation models consistently display stronger generalizability. Notably, Wave2Vec2BERT achieves the highest generalizability, which may be attributed to its large-scale and diverse pretraining data. Pretrained on 4.5 million hours of unlabeled audio in more than 143 languages, Wave2Vec2BERT benefits from both scale and diversity. This suggests that a well-designed self-supervised model trained on diverse speech data can extract general and discriminative features, making it more applicable across different datasets for audio deepfake detection. It is important to note that CLAP, unlike other speech foundation models, does not generalize well across datasets. This is likely due to its primary focus on environmental audio data during pretraining, resulting in the extraction of irrelevant features for speech audio. This observation underscores that not all foundation models are equally suited for audio deepfake detection tasks.
4.2.2. Results on SONAR dataset
@@
Table 4. Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
@@
4.2.3. Can generalizability increase with model size?
In Table 3, it can be observed that Whisper-large always outperforms Whisper-small across all three datasets. In particular, on the LibriSeVoc test set, Whisper-large achieves accuracy, AUROC, and EER of 0.9572, 0.9901, 4.279%, respectively, which improves by 2.27%, 0.64%, and 2.272%, than that of Whisper-small. This trend is more evident in the In-the-wild dataset, which is closer to real-world scenarios since this dataset consists of speech data sourced from the internet. Specifically, Whisper-large achieves accuracy, AUROC, and EER of 0.8848, 0.9552, and 11.518%, respectively, which improves by 6.381%, 5.27%, and 6.381%, than that of Whisper-small.
Table 5. Whisper model sizes.
@@ 4.2.3. Can generalizability increase with model size?
Building on the above observation, we extend our analysis with controlled experiments on the entire Whisper model family. Specifically, the Whisper family comprises five different model sizes: Whisper-tiny, Whisper-base, Whisper-small, Whisper-medium, and Whisper-large. Table 5 presents the number of model parameters of them. Specifically, each model is fine-tuned on the Wavefake training dataset using the same hyperparameters. Our results show that as model size increases, the generalizability of the models improves as well.
@@
Table 6. Generalization across existing audio deepfake datasets. All Whisper models are trained/finetuned on the Wavefake training set. Green and orange indicate the best and second-best performance, respectively.
@@
Table 7. Evaluation on SONAR dataset. Green and orange indicate the best and second-best performance, respectively.
@@
4.2.4. On the effectiveness and efficiency of few-shot fine-tuning for customized detection
Despite the challenges in generalizing across different datasets, we investigate whether there exist efficient solutions that can enhance models’ detection performance on those challenging subsets from SONAR dataset. To this end, we conduct a case study on Wave2Vec2BERT and HuBERT, as these models perform relatively poorly on the OpenAI and SeedTTS datasets but demonstrate competitive performance on other subsets. Specifically, we generate 100 additional fake audio samples using the OpenAI TTS API and randomly select another 100 fake audio samples from the SeedTTS test set for few-shot fine-tuning and fine-tune the models for 30 epochs with a learning rate of 0.00001 and a weight decay of 0.00005. Our study yields several interesting findings.
Figures 2(a) and 2(b) present the results of fine-tuning Wave2Vec2BERT and HuBERT using varying numbers of samples from OpenAI. Before fine-tuning, Wave2Vec2BERT and HuBERT only achieve accuracies of 0.7833 and 0.5658, respectively. Notably, with only 10 shots of fake speech data, Wave2Vec2BERT reaches an accuracy of approximately 0.97, while HuBERT’s accuracy increases significantly to approximately 0.85. Importantly, the models’ generalization to other datasets remains unchanged, demonstrating the effectiveness and efficiency of few-shot fine-tuning. However, as the number of fine-tuning samples increases, HuBERT’s test accuracy on the WaveFake test set shows a declining trend, which is also observed for Wave2Vec2BERT. This decrease in accuracy for some datasets as the number of fine-tuning examples increases can be attributed to the inherent challenge of catastrophic forgetting (McCloskey and Cohen, 1989) in deep learning models. This phenomenon can impact the model’s generalization ability and degrade its performance on certain datasets. In addition, with an increasing number of fine-tuning examples, the model may overfit to the specific characteristics of the fine-tuning dataset, especially if the dataset is small relative to the model’s capacity or lacks sufficient diversity.
Figure 2. Performance of few-shot fine-tuning for Wave2Vec2BERT and HuBERT with a varying number of few-shot audio samples from OpenAI and Seed-TTS, respectively. (a) Fine-tune Wave2Vec2BERT on OpenAI. (b) Fine-tune HuBERT on OpenAI; (c) Fine-tune Wave2Vec2BERT on Seed-TTS; and (d) Fine-tune HuBERT on Seed-TTS.
@@ 4.2.4. On the effectiveness and efficiency of few-shot fine-tuning for customized detection
4.2.5. How well can detection models generalize across languages?
To evaluate the cross-lingual generalization capabilities of detection models, we further conduct experiments on the MLAAD dataset (Müller et al., 2024), which encompasses 38 diverse languages. Table 8 presents the average performance metrics across all languages, with detailed language-specific results provided in Tables 1011, and 12. Our analysis reveals several significant findings regarding model generalization across languages.
Foundation models demonstrate remarkable cross-lingual generalization capabilities, despite being fine-tuned exclusively on English speech data. From Table 8, it can be observed that Wave2Vec2BERT achieves exceptional performance with an accuracy of 0.9901, AUROC of 0.9950, and EER of 0.9921%. Similarly, HuBERT and Wave2Vec2 also show strong performance, with accuracies of 0.9320 and 0.9139, respectively. This robust cross-lingual generalization may be attributed to the diverse multilingual pretraining data
Table 8. Generalization of different models across languages.
Model Accuracy AUROC EER (%)
LFCC-LCNN 0.6986 0.7661 30.1447
Spec.+ResNet 0.6001 0.6310 39.9947
RawNet2 0.4538 0.4379 54.6237
RawGATST 0.8061 0.8762 19.3921
AASIST 0.8461 0.9157 15.3868
CLAP 0.5136 0.5125 48.6395
Whisper-small 0.8276 0.9033 17.2395
Whisper-large 0.8325 0.9081 16.7474
Wave2Vec2 0.9139 0.9387 8.5947
HuBERT 0.9320 0.9745 6.7974
Wave2Vec2BERT 0.9901 0.9950 0.9921
these models are exposed to during their self-supervised learning phase. Their ability to learn language-agnostic speech representations that capture fundamental acoustic properties relevant to deepfake detection.
In contrast, traditional detection models show varying degrees of success in cross-lingual generalization. AASIST and RawGATST achieve respectable average accuracies of 0.846 and 0.806 on MLAAD, respectively. However, their performance significantly degrades on the SONAR dataset (accuracies of 0.6975 and 0.5939, as shown in Table 4(a)). The disparity in performance between MLAAD (containing primarily open-source TTS-generated audio) and SONAR provides crucial insights. Foundation models maintain relatively consistent performance across both datasets, while traditional detection models show significant degradation on SONAR. This observation suggests that currently the primary challenges in audio deepfake detection are more closely tied to the realism and quality of synthetic audio rather than language-specific characteristics.
4.2.6. On the generalization on fake audios generated by Voice Conversion
Table 9. Evaluation on Voice Conversion subset from ASVSpoof2019.
Model Accuracy($\uparrow$) AUROC($\uparrow$) EER (%)($\downarrow$)
LFCC-LCNN 0.623 0.667 37.44
ResNet_Spec. 0.660 0.714 34.05
RawNet2 0.430 0.391 58.72
RawGATST 0.671 0.738 32.86
AASIST 0.764 0.841 23.52
CLAP 0.494 0.495 50.30
Whisper-small 0.994 1.000 0.56
Whisper-large 0.931 0.978 6.93
Wave2Vec 0.883 0.952 11.64
HuBERT 0.971 0.994 2.87
Wave2VecBERT 0.974 0.995 2.58
In addition to evaluating TTS-generated audios, we assess the models’ detection performance on Voice Conversion (VC) technology using the VC subset from the ASVSpoof2019 dataset (Wang et al., 2020). The results, presented in Table 9, demonstrate that foundation models maintain strong generalization capabilities on VC tasks, while traditional detection models show significant performance degradation.
Among the foundation models, Whisper-small achieves near-perfect performance with an accuracy of 0.994, AUROC of 1.000, and EER(%) of 0.56. This is followed by Wave2VecBERT and HuBERT, which achieve accuracies of 0.974 and 0.971, AUROC of 0.995 and 0.994, and EERs(%) of 2.58 and 2.87, respectively. In contrast, traditional models like LFCC-LCNN and ResNet_Spec. struggle with limited generalization and high error rates. While better models such as AASIST show improvement over earlier approaches (accuracy: 0.764, AUROC: 0.841, EER(%): 23.52), they still lag significantly behind foundation models. CLAP’s poor performance (accuracy: 0.494, EER(%): 50.30) further emphasizes the importance of advanced architectures and diverse pretraining data for robust audio deepfake detection. These results highlight the superiority of foundation models in detecting VC-generated fake audio, demonstrating their potential for addressing the challenges posed by sophisticated audio manipulation techniques.
Table 10. Accuracy ($\uparrow$) of different detection models across languages.
Language LFCC ResNet Raw Raw AASIST CLAP Whisper Whisper Wave2Vec2 HuBERT Wave2Vec2
LCNN Spec. Net2 GATST small large BERT
Romanian 0.747 0.702 0.498 0.808 0.848 0.511 0.744 0.737 0.955 0.875 0.998
Croatian 0.960 0.823 0.403 0.973 0.970 0.424 0.946 0.973 0.999 0.977 0.999
Dutch 0.675 0.558 0.524 0.769 0.844 0.520 0.816 0.789 0.780 0.944 0.993
Latvian 0.579 0.761 0.417 0.979 0.976 0.537 0.906 0.921 1.000 0.981 1.000
Ukrainian 0.651 0.624 0.445 0.800 0.822 0.491 0.752 0.750 0.982 0.935 0.998
Irish 0.767 0.584 0.539 0.858 0.894 0.456 0.896 0.900 0.975 0.941 0.997
Polish 0.644 0.581 0.528 0.825 0.871 0.523 0.813 0.822 0.941 0.941 0.998
Lithuanian 0.780 0.723 0.582 0.944 0.970 0.468 0.756 0.732 1.000 0.956 0.999
Chinese 0.537 0.546 0.438 0.664 0.736 0.405 0.794 0.779 0.814 0.943 0.998
Greek 0.769 0.508 0.482 0.862 0.879 0.421 0.850 0.879 0.981 0.939 0.995
German 0.691 0.695 0.424 0.804 0.843 0.640 0.773 0.775 0.915 0.914 0.995
Turkish 0.758 0.578 0.432 0.779 0.820 0.562 0.867 0.840 0.809 0.952 0.996
Russian 0.631 0.540 0.386 0.646 0.728 0.567 0.719 0.701 0.941 0.915 0.992
Arabic 0.748 0.548 0.510 0.758 0.832 0.534 0.856 0.845 0.815 0.958 0.998
Spanish 0.606 0.548 0.440 0.772 0.860 0.542 0.687 0.690 0.930 0.915 0.995
Estonian 0.712 0.579 0.445 0.875 0.904 0.388 0.891 0.900 0.989 0.932 0.998
Thai 0.643 0.381 0.473 0.807 0.851 0.464 0.805 0.871 0.983 0.929 0.996
Bulgarian 0.705 0.583 0.462 0.849 0.892 0.505 0.867 0.877 0.994 0.936 0.998
Vietnamese 0.764 0.582 0.518 0.760 0.826 0.545 0.935 0.937 0.959 0.985 1.000
Maltese 0.777 0.563 0.482 0.871 0.904 0.411 0.896 0.910 0.990 0.937 0.997
Persian 0.480 0.621 0.327 0.819 0.713 0.712 0.899 0.914 0.991 0.939 0.998
English 0.547 0.663 0.478 0.683 0.740 0.445 0.711 0.675 0.599 0.891 0.776
Turkman 0.697 0.399 0.454 0.815 0.854 0.520 0.851 0.880 0.990 0.924 0.998
Hungarian 0.657 0.584 0.440 0.719 0.791 0.507 0.764 0.767 0.807 0.896 0.994
Swedish 0.709 0.593 0.337 0.827 0.904 0.451 0.889 0.890 0.983 0.934 0.997
Japanese 0.707 0.518 0.500 0.764 0.826 0.462 0.857 0.840 0.808 0.940 0.996
Bengali 0.589 0.594 0.523 0.606 0.632 0.518 0.556 0.640 0.690 0.729 0.958
Italian 0.514 0.565 0.488 0.806 0.843 0.508 0.874 0.874 0.968 0.943 0.998
Finnish 0.691 0.583 0.450 0.735 0.780 0.545 0.767 0.744 0.880 0.931 0.996
Hindi 0.727 0.578 0.491 0.768 0.838 0.528 0.857 0.877 0.817 0.957 0.998
Swahili 0.820 0.746 0.558 0.895 0.876 0.759 0.807 0.841 0.982 0.914 0.996
Slovak 0.724 0.605 0.345 0.886 0.907 0.453 0.893 0.898 0.994 0.938 0.997
Danish 0.791 0.626 0.352 0.886 0.907 0.593 0.897 0.912 0.989 0.938 0.998
French 0.652 0.575 0.400 0.796 0.838 0.598 0.764 0.764 0.971 0.916 0.996
Portuguese 0.730 0.599 0.467 0.749 0.807 0.562 0.844 0.841 0.788 0.957 0.996
Korean 0.748 0.546 0.464 0.761 0.836 0.536 0.866 0.888 0.882 0.953 0.998
Slovenian 0.874 0.765 0.291 0.951 0.944 0.408 0.917 0.927 0.995 0.965 0.996
Czech 0.744 0.635 0.450 0.762 0.847 0.498 0.867 0.836 0.847 0.947 0.997
Average 0.699 0.600 0.454 0.806 0.846 0.514 0.828 0.833 0.914 0.932 0.990
Table 11. AUROC ($\uparrow$) of different detection models across languages.
Language LFCC ResNet Raw Raw AASIST CLAP Whisper Whisper Wave2Vec HuBERT Wave2Vec
LCNN Spec. Net2 GATST small large BERT
Romanian 0.836 0.750 0.545 0.902 0.937 0.490 0.835 0.839 0.969 0.942 1.000
Croatian 0.994 0.892 0.340 0.998 0.997 0.377 0.987 0.994 1.000 0.997 1.000
Dutch 0.747 0.572 0.518 0.853 0.920 0.533 0.901 0.875 0.841 0.985 0.999
Latvian 0.627 0.818 0.372 0.999 0.998 0.521 0.970 0.970 1.000 0.996 1.000
Ukrainian 0.713 0.664 0.450 0.885 0.895 0.474 0.846 0.852 0.995 0.977 1.000
Irish 0.839 0.613 0.573 0.929 0.958 0.434 0.958 0.966 0.984 0.983 1.000
Polish 0.709 0.600 0.541 0.899 0.938 0.533 0.905 0.915 0.961 0.982 1.000
Lithuanian 0.869 0.773 0.565 0.991 0.997 0.430 0.842 0.812 1.000 0.991 1.000
Chinese 0.525 0.599 0.409 0.703 0.764 0.378 0.874 0.867 0.866 0.983 1.000
Greek 0.856 0.488 0.468 0.929 0.949 0.400 0.931 0.945 0.992 0.979 0.999
German 0.781 0.749 0.394 0.874 0.911 0.676 0.864 0.872 0.942 0.970 1.000
Turkish 0.844 0.637 0.388 0.861 0.911 0.596 0.943 0.919 0.868 0.989 1.000
Russian 0.664 0.570 0.334 0.713 0.809 0.571 0.807 0.798 0.957 0.968 1.000
Arabic 0.840 0.573 0.528 0.841 0.911 0.555 0.932 0.915 0.869 0.988 1.000
Spanish 0.656 0.586 0.415 0.854 0.934 0.535 0.788 0.787 0.948 0.962 0.999
Estonian 0.787 0.601 0.405 0.942 0.967 0.365 0.955 0.961 0.997 0.983 1.000
Thai 0.714 0.337 0.448 0.880 0.927 0.425 0.902 0.950 0.992 0.969 1.000
Bulgarian 0.772 0.579 0.450 0.925 0.957 0.493 0.942 0.949 1.000 0.986 1.000
Vietnamese 0.855 0.621 0.539 0.849 0.909 0.574 0.984 0.987 0.974 0.999 1.000
Maltese 0.859 0.575 0.490 0.941 0.966 0.397 0.959 0.978 0.998 0.986 1.000
Persian 0.484 0.706 0.223 0.880 0.811 0.740 0.958 0.975 0.997 0.964 1.000
English 0.573 0.725 0.472 0.735 0.810 0.418 0.786 0.763 0.650 0.933 0.825
Turkman 0.775 0.337 0.442 0.886 0.932 0.510 0.927 0.954 0.998 0.972 1.000
Hungarian 0.720 0.628 0.407 0.828 0.893 0.516 0.846 0.862 0.853 0.955 1.000
Swedish 0.773 0.593 0.330 0.896 0.956 0.434 0.955 0.955 0.992 0.982 1.000
Japanese 0.797 0.532 0.520 0.855 0.917 0.445 0.938 0.923 0.861 0.985 1.000
Bengali 0.643 0.637 0.522 0.630 0.681 0.553 0.632 0.741 0.770 0.820 0.991
Italian 0.542 0.598 0.470 0.891 0.917 0.503 0.949 0.949 0.978 0.981 1.000
Finnish 0.774 0.599 0.400 0.829 0.894 0.601 0.857 0.824 0.914 0.977 1.000
Hindi 0.807 0.594 0.504 0.847 0.920 0.544 0.939 0.953 0.875 0.989 1.000
Swahili 0.908 0.805 0.620 0.955 0.937 0.822 0.883 0.904 0.993 0.966 1.000
Slovak 0.811 0.617 0.309 0.947 0.967 0.436 0.961 0.971 0.999 0.985 1.000
Danish 0.873 0.667 0.307 0.955 0.954 0.593 0.961 0.977 0.997 0.984 1.000
French 0.713 0.615 0.361 0.884 0.916 0.638 0.835 0.849 0.978 0.967 0.999
Portuguese 0.821 0.650 0.460 0.831 0.899 0.560 0.924 0.904 0.851 0.989 1.000
Korean 0.835 0.557 0.469 0.843 0.920 0.543 0.941 0.955 0.915 0.989 1.000
Slovenian 0.948 0.846 0.224 0.990 0.992 0.354 0.972 0.977 1.000 0.991 1.000
Czech 0.828 0.676 0.428 0.849 0.929 0.505 0.940 0.919 0.899 0.990 1.000
Average 0.766 0.631 0.438 0.876 0.916 0.513 0.903 0.908 0.939 0.975 0.995
Table 12. EER (%) ($\downarrow$) of different detection models across languages.
Language LFCC ResNet Raw Raw AASIST CLAP Whisper Whisper Wave2Vec HuBERT Wave2Vec
LCNN Spec. Net2 GATST small large BERT
Romanian 25.30 29.80 50.20 19.20 15.20 48.90 25.60 26.30 4.50 1.25 0.20
Croatian 4.00 17.70 59.70 2.70 3.00 57.60 5.40 2.70 0.10 2.30 0.10
Dutch 32.50 44.20 47.60 23.10 15.60 48.00 18.40 21.10 21.90 5.60 0.70
Latvian 42.10 23.90 58.30 2.10 2.40 46.30 9.40 7.90 0.00 1.90 0.00
Ukrainian 34.90 37.60 55.50 20.00 17.80 50.90 24.80 25.00 1.80 6.50 0.20
Irish 23.30 41.60 46.10 14.20 10.60 54.40 10.40 10.00 2.50 5.90 0.30
Polish 35.60 41.90 47.20 17.50 12.90 47.70 18.70 17.80 5.90 5.90 0.20
Lithuanian 22.00 27.70 41.80 5.60 3.00 53.20 24.40 26.80 0.00 4.40 0.10
Chinese 46.30 45.40 56.20 33.60 26.40 59.50 20.60 22.10 18.60 5.70 0.20
Greek 23.10 49.20 51.80 13.80 12.10 57.90 15.00 12.10 1.90 6.10 0.50
German 30.90 30.50 57.60 19.60 15.70 36.00 22.70 22.50 8.50 8.60 0.50
Turkish 24.20 42.20 56.80 22.10 18.00 43.80 13.30 16.00 19.10 4.80 0.40
Russian 36.90 46.00 61.40 35.40 27.20 43.30 28.10 29.90 5.90 8.50 0.80
Arabic 25.20 45.20 49.00 24.20 16.80 46.60 14.40 15.50 18.50 4.20 0.20
Spanish 39.40 45.20 56.00 22.80 14.00 45.80 31.30 31.00 7.00 8.50 0.50
Estonian 28.80 42.10 55.50 12.50 9.60 61.20 10.90 10.00 1.10 6.80 0.20
Thai 35.70 61.90 52.70 19.30 14.90 53.60 19.50 12.90 1.70 7.10 0.40
Bulgarian 29.50 41.70 53.80 15.10 10.80 49.50 13.30 12.30 0.60 6.40 0.20
Vietnamese 23.60 41.80 48.20 24.00 17.40 45.50 6.50 6.30 4.10 1.50 0.00
Maltese 22.30 43.70 51.80 12.90 9.60 58.90 10.40 9.00 1.00 6.30 0.30
Persian 52.00 37.90 67.30 18.10 28.70 28.80 10.10 8.60 0.90 6.10 0.20
English 45.30 33.70 52.20 31.70 26.00 55.50 28.90 32.50 40.10 10.90 2.24
Turkman 30.30 60.10 54.60 18.50 14.60 48.00 14.90 12.00 1.00 7.60 0.20
Hungarian 34.30 41.60 56.00 28.10 20.90 49.30 23.60 23.30 19.30 10.40 0.60
Swedish 29.10 40.70 66.30 17.30 9.60 54.90 11.10 11.00 1.70 6.60 0.30
Japanese 29.30 48.20 50.00 23.60 17.40 53.80 14.30 16.00 19.20 6.00 0.40
Bengali 41.10 40.60 47.70 39.40 36.80 48.20 44.40 36.00 30.90 27.10 4.20
Italian 48.60 43.50 51.20 19.40 15.70 49.20 12.60 12.60 3.20 5.70 0.20
Finnish 30.90 41.70 55.00 26.50 22.00 45.50 23.30 25.60 12.00 6.90 0.40
Hindi 27.30 42.20 50.90 23.20 16.20 47.20 14.30 12.30 18.30 4.30 0.20
Swahili 18.00 25.40 44.20 10.50 12.40 24.10 19.30 15.90 1.80 8.60 0.40
Slovak 27.60 39.50 65.50 11.40 9.30 54.70 10.70 10.20 0.60 6.20 0.30
Danish 20.90 37.40 64.80 11.40 9.30 40.70 10.30 8.80 1.10 6.20 0.20
French 34.80 42.50 60.00 20.40 16.20 40.20 23.60 23.60 2.90 8.40 0.40
Portuguese 27.00 40.10 53.30 25.10 19.30 43.80 15.60 15.90 21.30 4.30 0.40
Korean 25.20 45.40 53.60 23.90 16.40 46.40 13.40 11.20 11.80 4.70 0.20
Slovenian 12.60 23.50 70.90 4.90 5.60 59.20 8.30 7.30 0.50 3.50 0.40
Czech 25.60 36.50 55.00 23.80 15.30 50.20 13.30 16.40 15.30 5.30 0.30
Average 30.14 40.00 54.62 19.39 15.39 48.64 17.24 16.75 8.59 6.80 0.99
5. Discussion